# Caution for the Environment: Multimodal Agents are Susceptible to Environmental Distractions

### v1

@@ 3.3 Measurement
The evaluation of the predicted action $\hat{a}$ is formalized separately for two kinds of agents, denoted as $\textsc{M}_{txt}$ and $\textsc{M}_{loc}$ in Eq. 5. (i) Strong generalist MLLMs (e.g., GPT-4o) can operate GUIs as agents with natural language by predicting actions with screen element descriptions, like “Submit button”. The prediction is measured by token-level $F_{1}$ with annotated actions and matches one if $F_{1}$ surpasses a threshold, $\tau_{txt}$. (ii) specialist agents (e.g., CogAgent) are trained to generate operating locations using precise coordinates on the screen. The predicted coordinate matches an annotated action if it falls into an annotated box,
@@
where the outputs are bool indicators. Based on the type of the matched labels, accuracy for gold actions, distracted actions, or invalid actions are computed respectively,
$\begin{split}&Acc_{gold}=\frac{1}{|D|}\sum_{d\in D}\exists a_{i}\in\{a_{gold}\},\textsc{M}(\hat{a},a_{i}),\\ &Acc_{dist}=\frac{1}{|D|}\sum_{d\in D}\exists a_{i}\in\{a_{dist}\},\textsc{M}(\hat{a},a_{i}),\\ &Acc_{inv}=1-\frac{1}{|D|}\sum_{d\in D}\exists a_{i}\in A,\textsc{M}(\hat{a},a_{i}),\\ \end{split}$ (6)
where $Acc_{\texttt{gold}}$ reflects the faithfulness and helpfulness of agents. $Acc_{\texttt{dist}}$ shows the unfaithfulness, i.e. how often agents are distracted from their goals. $Acc_{\texttt{inv}}$ indicates the multimodal capabilities including perception and prediction, i.e., how often agents fail to give valid actions.
3.4 Working Patterns
The behavior of agents can be sensitive to working patterns Shinn et al. (2024); Khattab et al. (2022), especially the perception level for complex environments like GUIs. For a comprehensive study, we implement working patterns with three levels of environmental perception, which are an implicitly-perceived level, a partially-perceived level, and a well-perceived level.
@@ 3.4 Working Patterns
$\circ$ CoT prompt. Chain-of-Thought (CoT) Wei et al. (2023a) have unveiled the reasoning capability of agents by asking the models to generate explicit rationales for deriving an answer Yao et al. (2022b); Hong et al. (2023). With a CoT-like pattern, the agent first receives the screenshot state to extract possible actions (“thoughts”), then predicts the next action based on the goal, denoted as
@@
$\circ$ Action annotations. Extracting all available actions from a screenshot is still a bottleneck for agents. If this burden is removed, the performance can depend more on judging distractions and keeping faithfulness. The available actions can be integrated into the input, denoted as
@@
In essence, providing available actions means two changes, as summarized in Table 2, (i) information for potential actions entailed in the image is disclosed and perceived as texts; (ii) information is fused into the text channel from the vision channel  [^1] Due to the space limitation, the prompts and detailed implementation will be released with our code and data..
@@ 4 Experiments
This section will introduce the implementation settings including the dataset and agent models and then show our empirical results with findings.
@@ 4.1 Implementation
Agent models. We implement a series of well-known MLLMs on our datasets. (i) Generalist agents. Multi-modal versions of strong black-box LLMs have shown promising performance and are available by API services, including GPT-4v, GPT-4o, GLM-4v GLM et al. (2024), Qwen-VL-plus Bai et al. (2023), and Claude-Sonnet-3.5Templeton et al. (2024). We also consider powerful open-source MLLMs, including Qwen-VL-chat Bai et al. (2023), MiniCPM-Llama3-v2.5 Hu et al. (2024), LLaVa-v1.6-34B Liu et al. (2023). (ii) Specialist agents. Recent studies released expert MLLMs for GUI agents after post-pre-training or instruction fine-tuning, including CogAgent-chat Hong et al. (2023) and SeeClick Cheng et al. (2024).
@@
Table 4: Results on the scenario of pop-up boxes.
@@
Table 5: Results on the scenario of search.
@@
Table 6: Results on the scenario of recommendation.
@@
Table 7: Results on the scenario of chat.
@@ 4.2 Main Results
Experimental results are shown in Table 3-7. Specifically, Table 3 shows an overview of the average of our four scenarios by direct prompt, and the following four tables present detailed scores across different scenarios and working patterns. Our results answer the following three key questions.
@@
(i) Can the multimodal environment distract a GUI agent from its goal? In risky environments, multimodal agents are susceptible to distractions that may lead them to abandon their goals and act unfaithfully. Each model produces actions that deviate from the original goal across our four scenarios. Such distracted predictions hinder the accuracy of gold actions. The Strong APIs (9.09% of GPT-4o) and specialist agents (6.84% of SeeClick) are more faithful than generalist open-source agents.
@@
(ii) What is the relation between faithfulness ($Acc_{\texttt{dist}}$) and helpfulness ($Acc_{\texttt{gold}}$)? There are two situations. First, MLLMs with strong capabilities can be both helpful and faithful ( GPT-4o, GPT-4v, and Claude). They exhibit low $Acc_{\texttt{inv}}$ scores, and relatively higher $Acc_{\texttt{acc}}$ and lower $Acc_{\texttt{dist}}$ (e.g., GPT-4o on pop-up boxes, search, and recommendation). Whereas, stronger perception but inadequate faithfulness can lead to greater susceptibility to distractions and lower helpfulness. For instance, GLM-4v demonstrates a higher $Acc_{\texttt{dist}}$ and a much lower $Acc_{\texttt{inv}}$ compared to open-sourced MLLMs, and GPT-4v and GPT-4o exhibit this trend in the chat subset. Hence, faithfulness and helpfulness are not mutually exclusive but can be enhanced simultaneously. It is even more critical to enhance faithfulness for stronger MLLMs.
@@
(iii) Can multimodal environmental perception help alleviate unfaithfulness? By implementing different working patterns, visual information is integrated into the textual channel to augment environmental perception. However, the results indicate that textual augmentation for GUI comprehensive can actually increase distractions. The increase of distraction action can outweigh the benefits, as seen in almost all setups with action annotations in the pop-up box, recommendation, and chat subsets (e.g., Qwen-VL, LLaVa, and GLM-4v). CoT-prompt, as a self-guided textual augmentation, can largely alleviate the perception burden but also increase distractions. These working patterns cannot work as a positive “defense” of environmental distractions.
This finding highlights two key points: firstly, there is a risk associated with stronger perception capabilities, and secondly, the fusion of UI information across textual and visual modalities (such as OCR) must be approached with greater caution.
@@ 5 Analysis
5.1 Comparing MLLMs
Among the generalist agents, GPT-4o demonstrates the best faithfulness and effectiveness in our scenarios, with the minimum average $Acc_{\texttt{distract}}$ (9.09%), and the maximum average $Acc_{\texttt{gold}}$ (74.3The open-sourced models get close scores on average, where LLaVa and MiniCPM are generally better. However, they demonstrate different abilities across scenarios. LLaVa is better at Searching and Recommending subsets, indicating advanced textual perception. MiniCPM is better at the pop-up boxes, and thus can be superior for visual (layouts or icons) knowledge.
Regarding specialist agents, the $Acc_{\texttt{dist}}$ of both CogAgent and SeeClick is much lower than general MLLMs, indicating that they enjoy higher faithfulness. CogAgent outperforms all agents except GPT-4 and Claude on both faithfulness and effectiveness. Shortcuts hinder the full potential of SeeClick, causing a high proportion of invalid actions. Specifically, when SeeClick encounters irrelevant pop-up boxes, it often predicts the coordinates at the very top right corner, likely attempting to close the box, rather than the actual position of the cross mark. And on screenshots of search pages, it often clicks the search bar. Once the available action annotations are input, the invalid actions and distractions are significantly reduced. This suggests that SeeClick knows what to do but draws wrong conclusions about the locations. These shortcuts can result from the inherited GUI knowledge gained during domain pre-training.
5.2 Comparing Working Patterns
Our three considered working patterns provide different levels of hints for the action prediction task. The direct pattern represents the environment with only an image. The action annotations expose the ground truth action space that could nearly substitute the environmental perception, making the task akin to a multiple-choice problem. This represents the upper bound of the perception capability. As a transition in between, CoT is applied to first ask the agent to predict a pseudo-action space, which is used to guide its action.
Our results show that the proportions of both gold actions and distracted actions largely increased with ground truth action space. However, on the other hand, the distracted proportions mean that even with a “perfect” perception, the agents are still vulnerable to distractions. For most models, the CoT prompt can work to provide some guidance and restrain agents’ behavior from invalid actions, but the distracted proportions also increased. Although it can not completely defend, the self-guided step-by-step process demonstrates the potential for mitigation.
5.3 Comparing Scenarios
The four simulated scenarios vary in emphasis and difficulty based on our empirical results. Figure 3 characters the variances in two types of challenges.
(ii) perception. In our results, the distractions are more successful in the recommendation subset. The chatting subset suffers from invalid actions or valid but unrelated actions. Accordingly, we also qualitatively illustrate the type and level of the perception difficulty. The pop-up boxes and chatting page mainly require the comprehension of the layout and icons. For example, find the cross mark to close the pop-up box or find the icon most related to the goal. The chatting page is more complex and implicit. For textual perception, true search results are more compact and closely related to the query. In contradiction, the real products for recommendation are noisy, more realistic but less attractive than the fake ones.
Figure 3: Illustration of scenario features.
6 Towards the Adversarial Perspective
The study discusses the general problem of environments that are imperfect, noisy, or flawed can distract multimodal agents, while they are not abnormal or malicious. However, such distractions can be exploited for malicious purposes. This section considers an extreme situation with an adversarial role and shows the feasibility of an active attack to distract GUI agents from the environment, named environment injection.
6.1 Threat Model
The user communicates with a multimodal GUI agent. The attacker needs to mislead the agent by only altering the GUI environment. The attacker can eavesdrop on the messages from the user and get the goal. The attacker can also hack the related environment to change the action space. For example, it is possible to block the package from a host and change the contents of a website. The problem is denoted as
@@ 6.1 Threat Model
6.2 Method and Experiments
Without loss of generality, we verified the feasibility of environment injection on the pop-up box scenario. The box layout is limited to one button to accept and one to reject. The box contents are distractions, thus the gold action is to click the button reject or the cross mark and the bad action is to click the button accept.
We propose and implement a brief but effective method to rewrite the two buttons. (i) The button to accept is rewritten to be ambiguous, and reasonable for both the distraction and the true goal. We find a common operation for those two purposes. But the contents in the box clarify the function of the buttons. We leverage the findings that such contexts in the screenshots can be ignored. (ii) The button to reject is rewritten to emotionally charged language. Such leading emotions can sometimes be persuasive or even manipulative tactics to influence user decisions. The phenomenon is common in APPs, like “Cruelly Leave” for uninstalling.
Algorithm 1 Attacking sample generation
  Initialize: Web environment $s$, User’s goal $g$, Agent $A$, Maximum tries $t_{m}$, LLM
  Notions: Distracting goal $d$, Pop-up box $B$, action prediction $a$
  $t\leftarrow 0$
  for $t<t_{m}$ do
     $d\leftarrow\text{LLM}(g,s)$, $B\leftarrow\text{LLM}(d)$
[…]
     $B.\textit{button\_rej}\leftarrow\text{LLM}(d)$
     $s\leftarrow s+B$
     $a\leftarrow A(s,g)$
[…]
        Attack success
     end if
     $t\leftarrow t+1$
  end for
@@ 6.2 Method and Experiments
Figure 4: Examples of simulated data.
@@ 7 Conclusion
This paper investigates the faithfulness of multimodal GUI agents and exposes the impact of distractions from the environment. We introduce a novel research question where both the user and the agent are benign, and the environment is not malicious but risky for distracting contents. We simulate distractions in four scenarios with localized annotations and implement three working patterns with varying perception levels. A broad range of generalist agents and specialist agents are evaluated. The experimental results demonstrate that vulnerability to distractions significantly diminishes both faithfulness and helpfulness. Additionally, we propose an attacking method called “environment injection”, which exploits unfaithfulness for malicious purposes by altering distractions to include ambiguous or emotionally misleading content. More importantly, this paper calls for increased collective focus on the faithfulness of multimodal agents. We suggest future work including pre-training for faithfulness alignment, considering the correlation between environment contexts and instructions, forecasting the possible consequences of executing actions and introducing human interaction when necessary.

### v2

@@ 3.3 Measurement
The measurement of the predicted action $\hat{a}$ is defined separately for two kinds of agents in Eq. 5. (i) Generalist MLLMs (e.g., GPT-4o) predict the operations on GUIs with natural language by describing screen elements as operating targets, like the “Submit button”. It is measured by token-level $F_{1}$ and matched with one annotated action if $F_{1}$ surpasses a threshold, $\tau_{txt}$. (ii) Specialist agents (e.g., CogAgent) are trained to generate operating locations using precise coordinates of the screen. The predicted coordinate matches an annotated action if it falls into an annotated box,
@@
where $\textsc{M}_{txt}$ and $\textsc{M}_{loc}$ are bool indicators. Next, based on the action labels, accuracy for gold actions, distracted actions, or invalid actions are computed respectively, where $Acc_{\texttt{gold}}$ reflects the faithfulness and helpfulness of agents; $Acc_{\texttt{dist}}$ shows the unfaithfulness, i.e., how often agents are distracted from their goals; $Acc_{\texttt{inv}}$ indicates how often agents fail to give valid actions, reflecting the overall capabilities,
$\begin{split}&Acc_{gold}=1/|D|\sum_{d\in D}\exists a_{i}\in\{a_{gold}\},\textsc{M}(\hat{a},a_{i}),\\ &Acc_{dist}=1/|D|\sum_{d\in D}\exists a_{i}\in\{a_{dist}\},\textsc{M}(\hat{a},a_{i}),\\ &Acc_{inv}=1-1/|D|\sum_{d\in D}\exists a_{i}\in A,\textsc{M}(\hat{a},a_{i}).\\ \end{split}$ (6)
3.4 Working Pattern
Agents can be sensitive to working patterns Shinn et al. (2024), particularly in complex environments: extracting available actions from a screen remains a bottleneck for GUI agents. For a comprehensive study, we implement three working patterns that gradually alleviate perception challenges (Table 2).
@@ 3.4 Working Pattern
$\circ$ CoT prompt. Chain-of-Thought (CoT) Wei et al. (2023a) have unlocked the reasoning capability of agents by generating intermediate rationales for deriving an answer. With a CoT-like pattern, the agent first receives the screenshot to extract possible actions (“thoughts”), then predicts the next action based on the goal, denoted as
@@
$\circ$ Action annotations. If the perception burden is removed, the agent’s behavior can depend more on judging distractions and keeping faithfulness. The available actions can be integrated into the input, denoted as
@@
In essence, providing available actions means two changes, as summarized in Table 2, (i) the action spaces are disclosed like multiple-choice questions; (ii) information is fused into the text channel from the vision channel. Appendix B.1 shows the prompts for each working pattern.
@@ 4 Experiments
We present empirical studies, including implementation and experimental results with key findings.
@@ 4.1 Implementation
Agent models. We implement a series of well-known MLLMs on our datasets. (i) Generalist agents. Multimodal versions of strong black-box LLMs have shown promising performance and are available by API services, including GPT-4v, GPT-4o, GLM-4v GLM et al. (2024), Qwen-VL-plus Bai et al. (2023), and Claude-Sonnet-3.5 Templeton et al. (2024). We also consider powerful open-source MLLMs, including Qwen-VL-chat-7B Bai et al. (2023), MiniCPM-Llama3-v2.5 Hu et al. (2024a), LLaVa-v1.6-34B Liu et al. (2023). (ii) Specialist agents. Recent studies released expert MLLMs for GUI agents after post-pre-training or instruction fine-tuning, including CogAgent-chat Hong et al. (2023) and SeeClick Cheng et al. (2024). Details are shown in Appendix B.
@@
Table 4: Results on the Pop-up box subset.
@@
Table 5: Results on the Search subset.
@@
Table 6: Results on the Recommendation subset.
@@
Table 7: Results on the Chat subset.
@@ 4.2 Main Results
Experimental results are shown in Table 3-7. Specifically, Table 3 shows an overview of the average of our four subsets with direct prompt, and the following four tables present detailed scores across different scenarios and working patterns. Our results answer the following three key questions.
@@
(i) Can the multimodal environment distract a GUI agent from its goal? Multimodal agents are susceptible to distractions that may lead them to abandon their goals and act unfaithfully. Each model produces actions that deviate from the original goal across our four scenarios. Such distracted predictions hinder the accuracy of gold actions. Strong APIs (9.09% of GPT-4o) and specialist agents (6.84% of SeeClick) are more faithful than generalist open-source agents. We also found “shortcut” in SeeClick, which suggests that GUI-domain pre-training facilitates the agent’s faithfulness but can also introduce shortcut knowledge. Detailed discussions are presented in Appendix A.1.
@@
(ii) What is the relation between faithfulness ($Acc_{\texttt{dist}}$) and helpfulness ($Acc_{\texttt{gold}}$)? There are two situations. First, MLLMs with strong overall capabilities can be both helpful and faithful (GPT-4o, GPT-4v, and Claude). They exhibit low $Acc_{\texttt{inv}}$ scores, and relatively higher $Acc_{\texttt{acc}}$ and lower $Acc_{\texttt{dist}}$ (e.g., GPT-4o on Pop-up box, Search, and Recommendation subsets). Whereas, stronger perception capability but inadequate faithfulness can lead to greater susceptibility to distractions and lower helpfulness. For instance, GLM-4v demonstrates a higher $Acc_{\texttt{dist}}$ and a much lower $Acc_{\texttt{inv}}$ compared to open-sourced MLLMs, because it successfully finds available actions but fails to decide on the correct one. GPT-4v and GPT-4o exhibit this trend in the Chat subset. Therefore, faithfulness and helpfulness are not mutually exclusive but can be enhanced simultaneously. It is even more critical to enhance faithfulness for stronger MLLMs.
@@
(iii) If we reduce the burden of environment perception by providing candidate actions, does the threat of environmental distractions still exist? By implementing different working patterns, visual information is integrated into the textual channel to augment environmental perception. However, the results indicate that textual prompts for candidate actions can not alleviate unfaithfulness and sometimes increase this risk. The increase of distracted action can outweigh the benefits, as seen in almost all setups with action annotations in the Pop-up box, Recommendation, and Chat subsets (e.g., Qwen-VL, LLaVa, and GLM-4v). CoT-prompt, as a self-guided textual augmentation, can largely alleviate the perception burden but also increase distractions. This finding highlights two key points: firstly, this unfaithfulness is associated with stronger perception, and secondly, the channel fusion across textual and visual modalities (such as OCR) must be approached with greater caution. More detailed analyses are in Appendices A.2 and A.3, including language-centric reasoning, specific phenomena, and subset comparison.
We summarize the two main challenges of environmental distractions as follows. The work of GUI agents is divided into environment understanding (perceiving) and decision-making for action (deciding). When perceiving, distractions cause significant changes in the action spaces. Pop-up boxes cover the screen with irrelevant content and disable appropriate actions. The chat record draws attention to a false action. When deciding, distractions also lead to inconsistency between the goal and the environmental contexts. This is similar to conflicts in the inputs, where LLMs can be misled by unexpected content Mallen et al. (2023); Wei et al. (2023b); Shi et al. (2023); Li et al. (2023).
@@ 5 Analysis
5.1 Towards Adversarial Perspective
Those distractions not only exist naturally in realistic environments, but also can be exploited for malicious purposes (Appendix C.2). This section considers the adversarial perspective and shows the feasibility of an active attack to mislead GUI agents, named environment injection.
5.1.1 Threat Model
The user communicates with a multimodal GUI agent. The attacker aims to mislead the agent by only altering the GUI environment. The attacker can eavesdrop on the messages from the user and reach their goal. The attacker can also hack the related environment to change the action space. For example, it is possible to block the package from a host and change the HTML contents, like man-in-the-middle. The problem is denoted as
@@ 5.1.1 Threat Model
5.1.2 Feasibility of Environment Injection
We verified the feasibility of environment injection on the pop-up box scenario. The box layout is simplified to one button to accept and one to reject. The box contents are distractions. Therefore, the gold action is to click the reject button or the cross mark, while the bad action is to accept.
We implement a brief but effective method to rewrite the pop-up box. (i) The button to accept is rewritten to be ambiguous, and reasonable for both the distraction and the true goal. Although the contents in the box clarify the actual function of the buttons, we found that agents often ignore contexts on the screen. (ii) The button to reject is rewritten to emotionally charged language. Such leading emotions can sometimes be persuasive or even manipulative tactics to influence user decisions. The phenomenon is common in APPs, like “Cruelly Leave” for uninstalling.
Different from Section 3.2, our attacker now has access to the user’s goal when writing distraction. Therefore, instead of Line 6 and Line 8 in Algo. 1, the adversarial distraction can be denoted to
$\begin{split}&d\leftarrow\text{LLM}(g,s),\\ &\textit{button\_acc}\leftarrow\text{LLM}(g,d),\\ &\textit{button\_rej}\leftarrow\text{LLM}(d)\end{split}$ (11)
@@ 5.1.2 Feasibility of Environment Injection
5.2 Towards the Faithfulness Improvement
Finally, we discuss the strategies to improve faithfulness against environmental distractions. Between the two summarized challenges above, we focus on the inconsistency of inputs, since the perception level has been discussed in different working patterns. We leave further study on the modality preference and alignment training strategy for future work.
5.2.1 Method
Differentiating the channel preference is a solution when dealing with inputs containing different information channels Lu et al. (2024); Wallace et al. (2024). We add a special token to distinguish the user’s goal from the environmental feedback and inject this preference by Direct Preference Optimization (DPO) Rafailov et al. (2024) training on a pseudo-dataset. Each data point includes several parallel inputs sampling from Alpaca Peng et al. (2023). By DPO, the model is trained to respond to the input tagged by the special token. Details are shown in Appendix B.2.
5.2.2 Experiments
This experiment trains Llama-3.1-8B-Instruct using LoRA Hu et al. (2022) on the pseudo-training set and tests on our Popup-box and Chat subsets following the Action Annotation working pattern. We compare the trained model after DPO with the baseline and original models with preference-aware prompts in Table 9.
Popup-box Chat
$Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$
Baseline 37.0 54.3 31.8 61.8
Prompt 33.3 51.0 24.5 70.9
DPO 37.3 55.7 40.9 53.6
Table 9: Results after DPO training.
After DPO, the user’s goal is highlighted and the performance on the Chat subset is improved significantly, while the improvement on the Popup-box subset is modest. The possible reason is that Popup-box subset requires excluding wrong actions rather than associating the user’s goal with the gold action, since the semantic distance between the gold action (rejecting the popup-box) and the user’s goal is relatively far. Moreover, Appendix C.3 suggests further improvement directions, i.e., visual-semantic reward and self-correction.
@@ 6 Conclusion
This paper investigates the faithfulness of multimodal LLM agents and exposes the impact of environmental distractions in GUI-based settings. We introduce a novel research question where both the user and the agent act benignly, while the environment, though non-malicious, contains distracting elements. To explore this, we simulate distractions and implement three working patterns with varying perception levels. A broad spectrum of generalist or specialist agents is evaluated. Our experimental results show that susceptibility to distractions significantly undermines both the faithfulness and helpfulness of agents. The core challenges of unfaithfulness are attributed to difficulties in perceiving the environment and making decisions based on inconsistent contexts. We further analyze the adversarial impacts of distractions and strategies to improve faithfulness. Overall, this work underscores the urgent need for increased research attention to the faithfulness of agents, particularly in preparation for their deployment in the real world.
Limitations
We acknowledge the limitations of this work. (i) We leave further explorations to improve the faithfulness for future work. Future efforts may explore pre-training with faithfulness-aligned objectives, modeling the correlation between environmental contexts and instructions, predicting the potential consequences of actions, and incorporating human feedback when necessary. Related analysis is provided in Section C.3. (ii) While our dataset simulates realistic distractions using a large language model, it does not cover all possible vulnerable scenarios. Future work could focus on constructing a more comprehensive set of distraction examples, potentially by leveraging crowd-sourced annotation of real-world websites. Further discussion and clarification are presented in Section C.1. (iii) This study quantifies LLM agent unfaithfulness based on objective accuracy across different action types. Investigating the severity of the resulting consequences is left for future work.